### v1

# Revisiting Monocular 3D Object Detection from Scene-Level Depth Retargeting to Instance-Level Spatial Refinement

Abstract
Monocular 3D object detection is challenging due to the lack of accurate depth. However, existing depth-assisted solutions still exhibit inferior performance, whose reason is universally acknowledged as the unsatisfactory accuracy of monocular depth estimation models. In this paper, we revisit monocular 3D object detection from the depth perspective and formulate an additional issue as the limited 3D structure-aware capability of existing depth representations (e.g., depth one-hot encoding or depth distribution). To address this issue, we propose a novel depth-adapted monocular 3D object detection network, termed RD3D, that mainly comprises a Scene-Level Depth Retargeting (SDR) module and an Instance-Level Spatial Refinement (ISR) module. The former incorporates the scene-level perception of 3D structures, retargeting traditional depth representations to a new formulation: Depth Thickness Field. The latter refines the voxel spatial representation with the guidance of instances, eliminating the ambiguity of 3D occupation and thus improving detection accuracy. Extensive experiments on the KITTI and Waymo datasets demonstrate our superiority to existing state-of-the-art (SoTA) methods and the universality when equipped with different depth estimation models. The code will be available.
1 Introduction
Input RGB Image
Figure 1: (a), (b), and (c) show VOXEL features from three different depth representations. (c) provides a more comprehensive and clearer scene structure feature. The last image surrounded by a dashed box represents instance-aware BEV features derived from (c), focusing on instance structure.
3D object detection is critical in computer vision and has a wide range of applications, such as autonomous driving, video surveillance, etc. Early LiDAR-based methods [40, 14, 32] have achieved promising performance but suffer from high hardware costs. Although several exclusively multi-view methods [28, 8, 17, 15] have gained considerable success, they typically require multiple cameras and suffer from view inconsistency. Recently, there have been considerable advancements in monocular methods [27, 21, 19] and garnered substantial attention from both academic and industrial communities, which give credit to their low cost and ease of deployment.
Monocular 3D object detection can be directly assisted by monocular depth estimation. However, inferring depth from a monocular image is inherently an ill-posed problem that affects the performance of monocular 3D object detection. Early depth-assisted methods [4, 37] resort to pre-training an auxiliary depth estimator, e.g., MonoDepth2 [7], to predict depth maps, but cannot achieve satisfactory detection performance. Most solutions (e.g., DD3D [24]) attribute it to the limited accuracy of monocular depth estimation models while overlooking the significance of depth representations. Recently, benefiting from the large-scale dataset and self-supervised training strategy, monocular depth estimation using large models (e.g., Depth Anything V2 (DAM v2) [41] and Metric3D v2 [42]) exhibit impressive accuracy and generalization. An intuitive idea is to leverage them to improve monocular 3D object detection. Unfortunately, as reported in Sec. 4.4, the results remain unsatisfactory, which drives us to rethink the well-recognized issue (i.e., limited depth accuracy problem) and revisit this type of approaches. In this work, we hold a different idea and contend that this issue also stems from the limited 3D structure awareness of existing depth representations (e.g., depth one-hot encoding or depth distribution).
To address this issue, we propose RD3D, a novel depth-adapted monocular 3D detection network. Different from existing approaches [30, 27] that adapt Lift-Splat-Shoot (LSS) [28] framework and bird’s-eye-view (BEV) space, RD3D retargets traditional depth representations in a new formulation: Depth Thickness Field to adapt to 3D object detection. The depth thickness field can not only reduce the impact of erroneous depth estimation but also integrate the thickness property for each image pixel in the 3D space. It implicitly describes the occupation thickness along camera rays, producing sharper and more accurate features, as illustrated in Fig. 1. RD3D consists of two main components. First of all, a Scene-Level Depth Retargeting (SDR) module is presented to retarget traditional depth representations to the proposed depth thickness field. Compared with previous representations, it embeds clear 3D structure awareness, benefiting in distinguishing different instances and backgrounds. Then, we sample image features and depth thickness field to establish the voxel space with predefined grids. Second, an Instance-Level Spatial Refinement (ISR) module is proposed to refine the voxel space by raising the oversight of individual instances. The network can focus more on the location and volume of instances rather than the whole 3D scene, eliminating the ambiguity of 3D occupation and thereby improving detection accuracy.
The main contributions are summarized as follows.
  • We investigate an additional issue of depth-assisted methods as the limited 3D structure-aware capability of existing depth representations and present RD3D, a depth-adapted monocular 3D detection network to solve it.
  • To incorporate the scene-level perception of 3D structures, we propose a Scene-Level Depth Retargeting (SDR) module that retargets traditional depth representations to a new one: depth thickness field, yielding sharp spatial features beneficial for 3D detection.
  • To further improve accuracy, an Instance-Level Spatial Refinement (ISR) module is proposed to refine the voxel space and eliminate the ambiguity of 3D occupation.
  • Compared with SoTA solutions, the proposed RD3D demonstrates superior accuracy and generalizability.
Figure 2: The overall framework of our proposed RD3D. The input image is first sent to the backbone to extract the features, and the depth map is transformed to depth one-hot encoding in parallel. The Scene-Level Depth Retargeting (SDR) module implicitly learns 3D scene structure features, retargeting depth one-hot encoding to depth thickness field (Sec. 3.3). The Instance-Level Spatial Refinement (ISR) module refines the voxel features with the guidance of instances, eliminating the ambiguity of 3D occupation (Sec. 3.4). Finally, the detector is applied to predict the 3D bounding boxes.
2 Related Work
2.1 Monocular 3D Detection
In recent years, substantial progress has been made in monocular 3D detection. Nevertheless, inferring 3D information from a 2D image presents an ill-posed problem. Previous works [9, 26, 23, 30] have employed a single module for the supervised learning of dense depth maps. DD3D [24] utilizes additional extensive unlabeled data from DDAD15M for depth pre-training, resulting in substantial enhancements in the performance of monocular 3D detection. More recently, MonoCD [39] suggests leveraging the geometric relationships among various depth cues to attain complementary form enhancements. Recent progress in monocular 3D detection has incorporated instance attributes [46, 18], which helps to tackle the inherent difficulties of this ill-posed task more effectively. Our method focuses on retargeting the traditional depth representations to a new formulation: depth thickness field, thereby improving the accuracy of 3D detection.
2.2 Monocular Depth Estimation
Monocular depth estimation aims to generate a single depth value for every pixel in an image. Neural network-based methods [36, 7, 31] often regress continuous depth by aggregating color and geometric structure information from an image. However, these methods are susceptible to overfitting to a single scenario, a common issue in monocular depth estimation. Recent methods [1, 34] have effectively mitigated the problem by creating large-scale relative depth datasets to learn relative relations. DAMv2 [41] unleashes the power of large-scale unlabeled data through semantic segmentation-assisted supervision and data augmentation. Metric3Dv2 [42] introduces a canonical camera transformation to resolve metric depth ambiguity across different camera setups, greatly improving depth accuracy and generalization. In our work, we utilize the depth map generated by the depth estimation model as a depth prior for the 3D detection model, guiding the generation of the depth thickness field representation.
2.3 3D Spatial Representations
Representing features in 3D space, such as BEV grids, removes scale differences in the targets and makes it easier for the network to learn feature scale consistency. Many works [15, 17, 8, 30, 22] employ BEV representations and achieve great success. LSS [28] generates point clouds by learning a pixel-level depth distribution and projecting it to obtain BEV features. BEVDepth [15] supervises the depth distribution using a projection of the depth map from the point cloud, thus improving the 3D detection performance. CaDDN [30] first used BEV representations for monocular 3D detection, and our approach follows its basic architecture design. Additionally, occupancy as an explicit representation of 3D space has recently experienced a surge in research. TPVFormer [10] employs sparse 3D occupancy labels from LiDAR as the supervision to obtain 3D features. OccupancyM3D [27] introduces occupancy to monocular 3D detection, exploring feature encoding and representation in frustum and voxel space. Unlike the above methods, we concentrate on combining occupancy representation and 3D bounding box (bbox) to explore the mutual support of occupancy and monocular 3D detection tasks.
3 Methodology
Figure 3: Comparisons between different depth representations. The two purple curves indicate the confidence of the pixel on each depth bin, obtained from three different depth representations. The left curve represents a thick object, and the right one represents a thin object. The depth thickness field has the unique ability to represent the object thickness along the camera ray.
3.1 Framework Overview
As illustrated in Figure 2, our framework takes a 3-channel RGB image and a single-channel depth map (generated by a pre-trained model) as input and outputs the 3D bounding box parameters of the object in the image. Initially, the RGB image $\mathbf{I}\in\mathbb{R}^{W_{I}\times H_{I}\times 3}$ is passed through the backbone (DLA34 [43]) to extract image features, while the depth map $\mathbf{D}\in\mathbb{R}^{W_{I}\times H_{I}}$ is converted into depth one-hot encoding representation. Next, the extracted image features $\mathbf{F}\in\mathbb{R}^{W_{F}\times H_{F}\times C}$ and depth one-hot encoding $\Psi\in\mathbb{R}^{W_{F}\times H_{F}\times D}$ are fed into the SDR module for further processing. Specifically, we employ the depth one-hot encoding representation to transform the extracted features into frustum features format $\mathbf{G}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times C}$. Then the depth thickness field $\mathbf{D}_{thickness}\in\mathbb{R}^{W_{F}\times H_{F}\times D}$ is adaptively learned under the supervision of our proposed soft-extended one-hot encoding. This depth thickness field is then used to refine the original image features and produce new frustum features $\mathbf{\tilde{G}}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times C}$. Following this, grid sampling is applied to voxelize the 3D features. The voxel features, denoted as $\mathbf{V}\in\mathbb{R}^{X\times Y\times Z\times C}$, are subsequently fed into the ISR module. This module refines the voxel space representation by learning an occupancy map, which is guided by the enhanced occupancy labels. Finally, the refined voxel features $\mathbf{\tilde{V}}\in\mathbb{R}^{X\times Y\times Z\times C}$ are compressed into a bird’s-eye view (BEV) space and forwarded to the 3D detection head to output the results.
3.2 Depth Thickness Field
Existing 3D object detection methods [26, 39] incorporate depth information to enhance 3D perception. However, depth estimation and 3D object detection are different tasks. Inappropriate depth representations often fail to provide meaningful cues for accurate detection. In this paper, we propose a novel depth representation method, termed Depth Thickness Field, which is more effective for 3D object detection. To clarify our contributions, we first review two commonly used depth representation methods—depth one-hot encoding and depth distribution—before detailing our proposed approach.
Depth One-Hot Encoding. Given a pixel located at position $(u,v)$ in a depth map, its depth value $z$ can be converted to $z^{\prime}$ using linear-increasing discretization (LID) [35]. Then we employ $D$ depth bins to represent the corresponding depth one-hot encoding of that pixel, which can be expressed as:
$\Psi(u,v)$ $=\left[\phi_{1},\phi_{2},\cdots,\phi_{d},\cdots,\phi_{D}\right],$ (1)
where
$\phi_{d}=\left\{\begin{array}[]{c}1\quad if\,\,d=z^{\prime},\\ 0\quad if\,\,d\neq z^{\prime}.\\ \end{array}\right.$ (2)
By applying Eq. 1 to each pixel in the depth map, we can obtain the depth one-hot encoding of the entire map, denoted as $\Psi$. This representation is sensitive to depth variations, aiding in the distinction of different objects. However, it heavily depends on the accuracy of the depth map. Moreover, this method generates a narrow response on the surface of the 3D object, limiting its ability to fully capture the three-dimensional structure.
Depth Distribution. Similar to depth one-hot encoding representation, the depth distribution of a pixel in the depth map can be represented as follows:
$P=\left[p_{1},p_{2},\cdots,p_{d},\cdots,p_{D}\right],\$ (3)
where $p_{d}$ represents the probability that the depth value $z^{\prime}$ lies within the $d$-th depth bin, satisfying the following constraint:
$\sum_{d=1}^{D}{p_{d}}=1,\quad p_{d}\geqslant 0.\$ (4)
Compared to depth one-hot encoding, depth distribution incorporates uncertainty when representing the depth probability distribution with depth bins, reducing the reliance on precise depth estimation. However, it still fails to fully capture the 3D structure (e.g., thickness) due to the absence of implicit thickness-aware guidance.
Depth Thickness Field. Unlike the two depth representations discussed above, we propose a novel depth thickness field method for mapping the given depth map. Our approach can implicitly represent the thickness of objects (as shown in Fig. 4), thereby enabling meaningful depth cues for 3D object detection. The details are as follows.
Figure 4: Thickness indicated by Depth Thickness Field along the camera ray.
Given a pixel in a depth map, its depth thickness field $F$ can be expressed as:
$F=\left[f_{1},f_{2},\cdots,f_{d},\cdots,f_{D}\right],\$ (5)
where $0\leqslant f_{d}\leqslant 1$, represents the value of the $d$-th depth bin along the camera ray.
The depth thickness field is not directly derived from the depth map. Instead, it is adaptively regressed by the network. While the depth distribution is also regressed, its supervision depends on pixel-wise ground truth depth values, which results in a limited range of responses within the depth bins. In contrast, our depth thickness field regression is guided by a soft-extended depth one-hot encoding constraint (detailed in Sec. 3.3), producing a broader range of responses (as shown in Fig. 3) and enabling an adaptive representation of object thickness.
3.3 Scene-Level Depth Retargeting
(a) Occupancy map from sparse point clouds.
(b) Occupancy map from sparse point clouds and 3D bboxes.
Figure 5: An slice example of occupancy map. Purple boxes indicate positive occupancy samples, with darker shades representing those generated from point clouds and lighter shades from 3D bounding boxes. Grey boxes indicate negative occupancy samples. (b) is more complete than (a).
These methods [30, 27] based on the LSS framework usually use depth information to explicitly transform 2D image features into 3D frustum features. However, due to the imperfect depth representation, the generated frustum features are not sufficient to perceive the 3D structure, resulting in inaccurate detection results. To address the challenges, we design the Scene-Level Depth Retargeting (SDR) module to retarget the traditional depth representations to the depth thickness field, thereby obtaining more meaningful frustum features for the subsequent 3D detection procedure. Specifically, given an RGB image $\mathbf{I}\in\mathbb{R}^{W_{I}\times H_{I}\times 3}$ and its corresponding depth map $\mathbf{D}\in\mathbb{R}^{W_{I}\times H_{I}}$, we first use a pre-trained backbone to extract image features $\mathbf{F}\in\mathbb{R}^{W_{F}\times H_{F}\times C}$. Simultaneously, we downsample the depth map and apply the LID method to produce $D$ depth bins for each pixel in the depth map. Subsequently, we employ Eq. 1 to obtain the depth one-hot encoding representation $\Psi\in\mathbb{R}^{W_{F}\times H_{F}\times D}$ of the depth map.
Given a pixel located at the position $(u,v)$ in the feature map, we can get the corresponding frustum feature grid $\textbf{G}(u,v)$ as follows:
$\textbf{G}\left(u,v\right)=\Psi\left(u,v\right)\otimes\textbf{F}\left(u,v\right),\$ (6)
where $\textbf{G}\left(u,v\right)$ $\in\mathbb{R}^{D\times C}$; $\otimes$ indicates the outer products. Applying Eq. 6 to each pixel in the feature map, we can get the whole frustum features $\textbf{G}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times C}$. Then we use three 3D convolution layers followed by a sigmoid function to refine the frustum features and regress the depth thickness field $\textbf{D}_{thickness}\in\mathbb{R}^{W_{F}\times H_{F}\times D}$. Finally, we can produce meaningful and progressive frustum features $\tilde{\textbf{G}}$ as follows:
$\tilde{\textbf{G}}=\textbf{D}_{thickness}\otimes\textbf{F},\$ (7)
where $\tilde{\textbf{G}}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times C}$.
As previously mentioned, the depth thickness field is adaptively generated by the network. To ensure that it accurately reflects object thickness, we impose the following constraints. Firstly, we define a mask to extend the response range within the standard depth one-hot encoding:
$M=[m_{1},m_{2},\cdots,m_{d},\cdots,m_{D}],$ (8)
where
$m_{d}=\left\{\begin{array}[]{rl}1&\quad|d-z^{\prime}|>l\,or\,d=z^{\prime},\\ 0&\quad else.\\ \end{array}\right.$ (9)
where $l$ represents the extended response range; $z^{\prime}$ has been defined in Sec. 3.2. Then the soft-extended depth one-hot encoding $\tilde{\Psi}(u,v)$ of a pixel in the depth map can be denoted as follows:
$\tilde{\Psi}(u,v)=\Psi(u,v)\odot M,$ (10)
where $\odot$ indicates the element-wise products. Finally, we introduce the focal loss [20] to formulate the objective function of this part as follows:
$L_{thickness}=\frac{1}{W_{F}\cdot H_{F}}FL(\tilde{\Psi},\textbf{D}_{thickness}).\$ (11)
3.4 Instance-Level Spatial Refinement
Figure 6: An slice example of occupancy map. The 3D bbox do not precisely conform to the object’s shape, resulting in false positive occupancy samples as the dark purple points.
Method Reference $Test,AP_{3D}|R_{40}$ $Test,AP_{BEV}|R_{40}$ $Val,AP_{3D}|R_{40}$
Easy Moderate Hard Easy Moderate Hard Easy Moderate Hard
GUPNet [23] CVPR21 22.26 15.02 13.12 30.29 21.19 18.20 22.76 16.46 13.72
MonoFlex [45] CVPR21 19.94 13.89 12.07 28.23 19.75 16.89 19.94 13.89 12.07
CaDDN [30] CVPR21 19.17 13.41 11.46 27.94 18.91 17.19 23.57 16.31 13.84
MonoDTR [9] CVPR22 21.99 15.39 12.73 28.59 20.38 17.14 28.84 20.61 16.38
DEVIANT [13] ECCV22 21.88 14.46 11.89 29.65 20.44 17.43 24.63 16.54 14.52
DID-M3D [26] ECCV22 24.40 16.29 13.75 32.95 22.76 19.83 22.98 16.12 14.03
MonoDDE [16] CVPR22 24.93 17.14 15.10 33.58 23.46 20.37 26.66 19.75 16.72
MonoNeRD [38] ICCV23 22.75 17.13 15.63 31.13 23.46 20.97 22.03 15.44 13.99
MonoDETR [44] ICCV23 25.00 16.47 13.58 33.60 22.11 18.60 28.84 20.61 16.38
MonoCD [39] CVPR24 25.53 16.59 14.53 33.41 22.81 19.57 26.45 19.37 16.38
MonoLSS [18] 3DV24 26.11 19.15 16.94 34.89 25.95 22.59 25.91 18.29 15.94
MonoDiff [29] CVPR24 30.18 21.02 18.16 - - - 32.18 22.02 19.84
OccupancyM3D [27] CVPR24 25.55 17.02 14.79 35.38 24.18 21.37 26.87 19.96 17.15
MonoMAE [11] NeurIPS24 25.60 18.84 16.78 34.15 24.93 21.76 30.29 20.90 17.61
RD3D(Ours) - 32.02 21.19 18.80 42.67 28.92 25.89 36.17 26.76 23.67
Improvement v.s. second-best +1.84 +0.17 +0.64 +7.29 +2.97 +3.3 +3.99 +4.74 +3.83
Table 1: Comparisons on KITTI test and val sets for Car category. Results are shown using the $AP(IoU=0.7)|R_{40}$ metric. We indicate the highest result with red and the second highest with blue. The performance metrics for the other methods are reported from the respective published results.
OccupancyM3D [27] enhances 3D spatial features by learning an occupancy representation in voxel space. However, the labels used to guide the occupancy representation generation emanate from sparse point clouds, which are confined to the object’s surface (as shown in Fig. 5(a)). As a result, they fail to accurately capture the thickness along the camera ray. Additionally, the limited perception range of LiDAR further restricts the ability to impose effective constraints on distant objects, exacerbating the ill-posed problem in the labels. This limitation significantly impacts the network’s performance. To address this challenge, we design the Instance-Level Spatial Refinement (ISR) module to refine the voxel spatial representation with the guidance of instances.
Given voxel features $\textbf{V}\in\mathbb{R}^{X\times Y\times Z\times C}$, we use 3D convolution with a sigmoid function to obtain the voxel occupancy features $\textbf{O}\in\mathbb{R}^{X\times Y\times Z\times 1}$ which then are employed to refine the previous voxel features. This procedure can be expressed as follows:
O $=\text{Sigmoid}\left(f\left(\textbf{V}\right)\right),$ (12)
[…]
where $f(\cdot)$ denotes a 3D hourglass-like design [2] and $\odot$ denotes the element-wise products.
Occupancy Labels. To generate more adapted occupancy labels for 3D object detection, we propose leveraging 3D bounding boxes to assist sparse point clouds in representing object thickness, as illustrated in Figure 5(b). Similar to OccupancyM3D [27], we denote the 3D occupancy labels derived from sparse point clouds as $O_{point}^{*}\in\mathbb{R}^{X\times Y\times Z}$, where $X$, $Y$, and $Z$ are determined by the predefined voxel size and detection range.
Considering that the ISR module learns the occupancy representation within the voxel space and that the original 3D bounding boxes do not precisely conform to the object’s shape, directly using them would result in ambiguous occupancy labels. To address this issue, we propose constructing refined 3D bounding boxes, derived from the originals, as shown in Fig. 6. These adjusted boxes retain the same center but are scaled down to reduce ambiguity in the occupancy labels. Specifically, let $\tilde{\textbf{B}}\in\mathbb{R}^{N\times 7}$ represent the refined 3D bounding boxes. For each voxel block within the voxel space, if it falls within any of the 3D bounding boxes, we designate it as occupied. Using this approach, we obtain additional 3D occupancy labels $O_{bbox}^{*}\in\mathbb{R}^{X\times Y\times Z}$ derived from the 3D bounding boxes. The final occupancy label is then obtained by taking the union of $O_{point}^{*}$ and $O_{bbox}^{*}$, i.e., $O^{*}=O_{point}^{*}\cup O_{bbox}^{*}$.
The ISR module enhances the perception of instances by adding occupancy labels from the 3D bounding boxes and then using it to refine the voxel spatial representation which is beneficial to improve detection accuracy.
3.5 Loss Function
Our objective function consists of three components: the first for 3D object detection, the second for learning occupancy in 3D voxel space, and the third for learning the depth thickness field. The first component follows the approach of CaDDN [30], denoted as $\mathcal{L}_{org}$. For the occupancy component, we adopt the loss function from OccupancyM3D [27], represented as $\mathcal{L}_{occ}$. The third component involves the depth thickness field, for which we use the thickness loss $\mathcal{L}_{thickness}$, as defined in Section 3.3. The overall objective function is then formulated as follows:
$\mathcal{L}=\mathcal{L}_{org}+\mathcal{L}_{occ}+\mathcal{L}_{thickness}\$ (13)
4 Experiments
Method Reference $3DmAP/mAPH(IoU=0.7)$ $3DmAP/mAPH(IoU=0.5)$
Overall 0 - 30m 30 - 50m 50m - $\infty$ Overall 0 - 30m 30 - 50m 50m - $\infty$
LEVEL 1
CaDDN [30] CVPR21 5.03/4.99 14.54/14.43 1.47/1.45 0.10/0.10 17.54/17.31 45.00/44.46 9.24/9.11 0.64/0.62
DEVIANT [13] ECCV22 2.69/2.67 6.95/6.90 0.99/0.98 0.02/0.02 10.98/10.89 26.85/26.64 5.13/5.08 0.18/0.18
DID-M3D [26] ECCV22 -/- -/- -/- -/- 20.66/20.47 40.92/40.60 15.63/15.48 5.35/5.24
MonoNeRD [38] ICCV23 10.66/10.56 27.84/27.57 5.40/5.36 0.72/0.71 31.18/30.70 61.11/60.28 26.08/25.71 6.60/6.47
MonoLSS [18] 3DV24 3.71/3.69 9.82/9.75 1.14/1.13 0.16/0.16 13.49/13.38 33.64/33.39 6.45/6.40 1.29/1.26
MonoDiff [29] CVPR24 -/- -/- -/- -/- 32.28/31.49 63.94/62.13 25.91/25.47 7.51/7.34
OccupancyM3D [27] CVPR24 10.61/10.53 29.18/28.96 4.49/4.46 0.41/0.40 28.99/28.66 61.24/60.63 23.25/23.00 3.65/3.59
RD3D(Ours) - 18.04/17.79 47.57/46.96 9.09/9.00 0.64/0.63 42.06/41.24 80.81/79.46 40.42/39.81 6.41/6.21
LEVEL 2
CaDDN [30] CVPR21 4.49/4.45 14.50/14.38 1.42/1.41 0.09/0.09 16.51/16.28 44.87/44.33 8.99/8.86 0.58/0.55
DEVIANT [13] ECCV22 2.52/2.50 6.93/6.87 0.95/0.94 0.02/0.02 10.29/10.20 26.75/26.54 4.95/4.90 0.16/0.16
DID-M3D [26] ECCV22 -/- -/- -/- -/- 19.37/19.19 40.77/40.46 15.18/15.04 4.69/4.59
MonoNeRD [38] ICCV23 10.03/9.93 27.75/27.48 5.25/5.21 0.60/0.59 29.29/28.84 60.91/60.08 25.36/25.00 5.77/5.66
MonoLSS [18] 3DV24 3.27/3.25 9.79/9.73 1.11/1.10 0.15/0.15 13.12/13.02 33.56/33.32 6.28/6.22 1.15/1.13
MonoDiff [29] CVPR24 -/- -/- -/- -/- 30.73/30.48 63.86/62.92 25.28/24.86 6.43/6.29
OccupancyM3D [27] CVPR24 10.02/9.94 28.38/28.17 4.38/4.34 0.36/0.36 27.21/26.90 61.09/60.49 22.59/22.34 3.18/3.13
RD3D(Ours) - 16.77/16.54 46.94/46.34 8.76/8.67 0.55/0.54 39.20/38.44 79.93/78.60 39.04/38.45 5.56/5.39
Table 2: Comparisons on Waymo val set for Vehicle category. We indicate the highest result with red and the second highest with blue. The performance metrics for the other methods are reported from the respective published results.
Exp. SDR ISR $Test,AP_{3D}|R_{40}/AP_{BEV}|R_{40}$ $Val,AP_{3D}|R_{40}/AP_{BEV}|R_{40}$
Easy Moderate Hard Easy Moderate Hard
(a) 26.95/37.64 16.15/23.26 13.45/19.77 28.15/41.71 21.62/31.53 19.11/28.34
(b) 31.12/40.76 20.04/27.49 17.33/24.29 36.71/51.60 25.52/35.12 22.17/31.83
(c) 31.97/42.29 19.79/27.28 17.38/24.27 35.67/48.06 24.96/33.60 21.50/29.49
(d) 32.02/42.67 21.19/28.92 18.80/25.92 36.17/47.42 26.76/35.68 23.67/31.94
Table 3: Main ablations on KITTI test and val sets. “Exp.” means experiment; “SDR” refers to Scene-Level Depth Retargeting module; “ISR” denotes Instance-Level Spatial Refinement module.
4.1 Implementation Details
We employ PyTorch [25] for implementation. The network is trained on 3 NVIDIA 4090 (24G) GPUs. We use Adam [12] optimizer with an initial learning rate of 0.001 and employ the one-cycle learning rate policy [33]. On KITTI dataset [6], we train the model for 80 epochs with a total batch size of 12 and use [2, 46.8] $\times$ [-30.08, 30.08] $\times$ [-3, 1] (meter) for the range. On Waymo dataset [5], we downsample the input size to 640 $\times$ 960 and train 12 epochs with a detection range [2, 59.6] $\times$ [-25.6, 25.6] $\times$ [-2, 2] (meter). The voxel size is [0.16, 0.16, 0.16] (meter).
4.2 Benchmarks and Metrics
KITTI 3D. KITTI 3D is the most widely used benchmark [6] for 3D object detection, consisting of 7,481 training samples and 7,518 testing samples. Only annotations of the training set are made public. The training set is commonly divided into KITTI train with 3712 samples and KITTI val with 3769 samples following [3]. There are three difficulty levels, easy, moderate, and hard, defined by bounding box height and occlusion level. Following common practice, we use $AP_{3D}|_{R_{40}}$ and $AP_{BEV}|_{R_{40}}$ under IoU threshold of 0.7 to evaluate the performance.
Waymo. The Waymo Open Dataset [5] is a large dataset that contains 798 sequences for training and 202 sequences for validation. Following previous works [30], we report the performance on Waymo val set for the vehicle category using samples only from the front camera. Results on Waymo val are measured by officially evaluation of the mean average precision (mAP) and the mean average precision weighted by heading (mAPH).
Exp. Depth Model SDR&ISR $Test,AP_{3D}|R_{40}/AP_{BEV}|R_{40}$ $Val,AP_{3D}|R_{40}/AP_{BEV}|R_{40}$
Easy Moderate Hard Easy Moderate Hard
(1) MonoDepth2* [7] 20.41/28.33 12.03/17.45 10.30/14.89 20.39/31.34 14.90/22.56 12.42/19.05
(2) MonoDepth2* 21.65/29.71 14.13/19.42 11.73/16.99 22.01/31.57 16.54/23.62 14.17/20.46
Improvement +1.24/+1.38 +2.1/+1.97 +1.43/+2.1 +1.62/+0.23 +1.64/+1.06 +1.75/+1.41
(5) DAMv2$\dagger$ [41] 26.90/38.51 15.40/22.86 12.48/19.26 35.24/48.36 20.97/29.94 17.09/24.65
(6) DAMv2$\dagger$ 32.74/41.72 20.19/27.11 17.04/23.36 39.82/52.30 27.46/37.79 24.12/33.74
Improvement +5.84/+3.21 +4.79/+4.25 +4.56/+4.10 +4.58/+3.94 +6.49/+7.85 +7.03/+9.09
(7) Metric3Dv2 [42] 26.95/37.64 16.15/23.26 13.45/19.77 28.15/41.71 21.62/31.53 19.11/28.34
(8) Metric3Dv2 32.02/42.67 21.19/28.92 18.80/25.89 36.17/47.42 26.76/35.68 23.67/31.94
Improvement +5.25/+5.03 +5.04/+5.66 +5.35/+6.12 +8.02/+5.71 +5.14/+4.15 +4.56/+3.60
Table 4: Ablations of different depth estimation models on KITTI test and val sets. * denotes the lightweight model and $\dagger$ indicates the further fine-tuned model. “DAMv2” refers to Depth Anything V2 [41].
Exp. Representation $Test,AP_{3D}|R_{40}/AP_{BEV}|R_{40}$
Easy Moderate Hard
1) One-Hot 26.95/37.64 16.15/23.26 13.45/19.77
2) Distribution 30.37/40.42 19.16/26.69 17.02/24.12
3) Thickness 32.02/42.67 21.19/28.92 18.80/25.89
Table 5: Ablation of depth representation. One-Hot: Depth One-Hot Encoding; Distribution: Depth Distribution; Thickness: Depth Thickness Field.
4.3 Main Results
KITTI 3D. Table 1 shows the main results of RD3D on KITTI test and val sets. Our method achieves the best results under three-level difficulty settings in both $AP_{BEV}$ and $AP_{3D}$. On the test set, RD3D exceeds all existing methods and surpasses the second-best under three-level difficulties by +1.84/+0.17/+0.64 in $AP_{3D}$, and by +7.29/+2.97/+3.33 in $AP_{BEV}$. In particular, RD3D does not require the depth ground truth (GT) which is hard to obtain for supervision when training the detection model. Guided by accurate depth priors, our method transforms traditional depth representations into a more effective depth thickness field for 3D detection, resulting in a significant improvement over existing approaches.
Waymo. We also evaluate our method on Waymo Open Dataset (Waymo) [5]. As shown in Table 2, our method obtains promising results, significantly outperforming existing methods in 0-30m, 30-50m, and Overall. For example, under LEVEL 1 setting, RD3D outperforms the second best by 7.38/7.23 mAP/mAPH (18.04/17.79 vs. 10.66/10.56) and 9.78/9.75 mAP/mAPH (42.06/41.24 vs. 32.28/31.49) with IoU 0.7 and 0.5 criterions respectively; under LEVEL 2 setting, RD3D outperforms the second best by 6.74/6.60 mAP/mAPH (16.77/16.54 vs. 10.03/9.94) and 8.47/7.96 mAP/mAPH (39.20/38.44 vs. 30.73/30.48) with IoU 0.7 and 0.5 criterions respectively. This success can be attributed to the effective depth thickness field representation. For objects within [50m, $\infty$], our method performs relatively worse. This is because our method is voxel-based, which has a detection range limitation: [2, 59.6] (meters).
4.4 Ablations
To illustrate the effectiveness of SDR and ISR modules proposed in this paper, we carry out ablations on both KITTI test and val sets as shown in Table 3. We compare the performance on Car category under IoU criterion 0.7.
SDR: Exp. (a) is the baseline which directly uses depth one-hot encoding to generate 3D features for 3D object detection. In Exp. (b), when retargeting depth representation from depth one-hot encoding to depth thickness field, the detection $AP_{3D}$ increases from 26.95/16.15/13.45 to 31.12/20.04/17.33 on the test set. ISR: As shown in Exp. (c), when enforcing the refinement of voxel spatial representation via learning 3D occupation with the guidance of instances, the detection $AP_{3D}$ is boosted to 31.97/19.79/17.38 on the test set. Finally, we can see RD3D obtains +5.25/+5.04/+5.35 in $AP_{3D}$ compared to the baseline from Exp. (d).
Universality Analysis. In Table 4, we report the performances of RD3D when equipped with different depth estimation models. DAMv2 and Metric3Dv2 are two large models for depth estimation, showing significant improvements in detection performance when integrated with our method. We have fine-tuned DAMv2 which was initially designed as a relative depth model on KITTI dataset for absolute depth estimation. Its results demonstrate that our method is well-suited to the current paradigm in depth estimation models: initial self-supervised training on relative depth, followed by fine-tuning for absolute depth. As an early lightweight model for depth estimation, Monodepth2 still benefits from our method for monocular 3D object detection. These experiments demonstrate both the effectiveness and the universality of RD3D.
vs. Traditional Depth Representations. In Table 5, we report the performance when replacing the depth thickness field with traditional depth representations. We can observe that the depth thickness field achieves better performance compared to depth one-hot encoding and depth distribution.
5 Conclusion
In this paper, we propose a depth-adapted network for monocular 3D object detection. Unlike current approaches, we reformulate a new key factor of the depth-assisted method from depth accuracy to the adaptability of depth representation and introduce a novel depth representation. The proposed SDR module retargets traditional depth representations to depth thickness field with self-supervision. We also introduce the ISR module to refine the voxel spatial representation with the guidance of instances. We conduct experiments on the challenging KITTI and Waymo open datasets. The results demonstrate our superiority to existing state-of-the-art (SoTA) methods and the universality when equipped with different depth estimation models.